# Koby Hayashi

### v1

@@ 2.1.3 Projected Gauss-Newton with Conjugate Gradients for SymNMF
The algorithm PGNCG-SymNMF was proposed for efficiently computing SymNMF in highly parallel computing environments [15]. It is an all-at-once method and uses the Projected Gauss-Newton method to directly optimize the SymNMF objective, Equation 2, and is an all-at-once method. The main computational load lies in solving a least squares problem of the form
@@
for a search direction ${\bm{\mathbf{{p}}}}$ at every iteration. The matrix ${\bm{\mathbf{{J}}}}$ and the vector ${\bm{\mathbf{{r}}}}$ are the Jacobian and residual of Equation 2 respectively. A solution to this LS problem is then approximated using the Conjugate Gradient (CG) method on the Normal Equations : ${\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{J}}}}{\bm{\mathbf{{p}}}}=-{\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{r}}}}$. A core computational kernel of the CG method is computing matrix vector products with the matrix ${\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{J}}}}$. Fortunately, the Jacobian ${\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{J}}}}$ has the Kronecker product form ${\bm{\mathbf{{J}}}}=-({\bm{\mathbf{{H}}}}\otimes{\bm{\mathbf{{I}}}}_{m})-{\bm{\mathbf{{P}}}}_{m,m}({\bm{\mathbf{{H}}}}\otimes{\bm{\mathbf{{I}}}}_{m})$, where ${\bm{\mathbf{{P}}}}_{m,m}$ is the perfect shuffle or “vec” permutation, which can be used to efficiently apply ${\bm{\mathbf{{J}}}}$ to a vector. Additionally the vector ${\bm{\mathbf{{g}}}}={\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{r}}}}$ has the form ${\bm{\mathbf{{g}}}}=-2\text{vec}({\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}-{\bm{\mathbf{{H}}}}({\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}))$, which is typically the main computational bottleneck and requires the matrix multiplications ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$. See [15] for details.
@@ 2.2.1 Randomized Range Finder
2:${\bm{\mathbf{{Q}}}}_{Y}\in\mathbb{R}^{m\times(r+\rho)}$ is an approximate orthonormal basis for the leading column span of ${\bm{\mathbf{{X}}}}$:
@@
Pseudocode for the the RRF is given in algorithm RRF. Parameters of the RRF are the target rank $r$, a column oversampling parameter $\rho$ (with $l=r+\rho$), and $q$, the number of power iterations to perform. The computational complexity of the RRF is $O(qmnl+ml^{2})$. The approximate output from the RRF is often used to compute a so called QB-Decomposition. That is if a matrix ${\bm{\mathbf{{X}}}}$ is input to the RRF and a matrix ${\bm{\mathbf{{Q}}}}_{X}$ is output then ${\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{Q}}}}_{X}^{\sf T}{\bm{\mathbf{{X}}}}={\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$, where ${\bm{\mathbf{{B}}}}_{X}={\bm{\mathbf{{Q}}}}_{X}^{\sf T}{\bm{\mathbf{{X}}}}$, is called a QB-Decomposition of ${\bm{\mathbf{{X}}}}$.
@@
For symmetric input, an approximate eigenvalue decomposition can also be obtained by using the approximate basis from the RRF. This procedure is shown in algorithm Apx-EVD, which stands for approximate eigenvalue decomposition. More details and references can be found in [18].
@@ 2.2.2 Leverage Score Sampling for Ordinary Least Squares
Another approach that we will use from RandNLA is sketching for ordinary least squares (OLS) problems, specifically, leverage score sampling for OLS problems. A standard OLS or $l2-$regression problem can be stated mathematically as
$\min_{{\bm{\mathbf{{x}}}}}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}$ (16)
@@
$\hat{{\bm{\mathbf{{x}}}}}_{ols}=\operatorname*{arg\,min}_{{\bm{\mathbf{{x}}}}}\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{S}}}}{\bm{\mathbf{{b}}}}\|_{2},$ (17)
where ${\bm{\mathbf{{S}}}}\in\mathbb{R}^{s\times m}$ with $s\ll m$ is called a sketching matrix. Computational savings come from the fact that one can now solve the smaller problem in Equation 17 as opposed to the full problem in Equation 16.
@@
$l_{i}({\bm{\mathbf{{A}}}})=\|{\bm{\mathbf{{Q}}}}_{A}[i,:]\|_{2}^{2}$ (18)
where the matrix ${\bm{\mathbf{{Q}}}}_{A}$ is any othonormal basis for the column space of ${\bm{\mathbf{{A}}}}$ and ${\bm{\mathbf{{Q}}}}_{A}[i,:]$ is the $i$th row of ${\bm{\mathbf{{Q}}}}_{A}$. For example the matrix ${\bm{\mathbf{{U}}}}_{A}$, where ${\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}^{\sf T}_{A}={\bm{\mathbf{{A}}}}$ is an thin SVD of ${\bm{\mathbf{{A}}}}$, can be used to calculate the leverage scores. These values are normalized into probabilities $p_{i}=\frac{l_{i}({\bm{\mathbf{{A}}}})}{\|{\bm{\mathbf{{Q}}}}_{A}\|_{F}^{2}}$. Using the $p_{i}$’s, $s$ samples are drawn with replacement and the matrix ${\bm{\mathbf{{S}}}}\in\mathbb{R}^{s\times m}$ is formed as
${\bm{\mathbf{{S}}}}_{ji}=\begin{cases}\frac{1}{\sqrt{sp_{i}}},\ \text{if row $i$ was drawn as the $j$th sample}\\ 0\end{cases}.$ (19)
@@ 3 NMF with Low-rank Approximate Input
Our first proposed algorithm is a method called Low-rank Approximate-Input NMF (SymNMF). LAI-NMF computes an NMF of a low-rank approximation of the initial data matrix ${\bm{\mathbf{{X}}}}$. The objective function for LAI-NMF is
$\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F},$ (20)
where ${\bm{\mathbf{{U}}}}_{X}\in\mathbb{R}^{m\times l}$, ${\bm{\mathbf{{V}}}}_{X}\in\mathbb{R}^{l\times n}$ with $k\leq l\ll\min(m,n)$ and ${\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}\approx{\bm{\mathbf{{X}}}}$ is a low-rank approximation of ${\bm{\mathbf{{X}}}}$. The primary idea is that an approximate solution to Equation 20 can be quickly computed by exploiting the product form of ${\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}$ to compute matrix vector products. That is, ${\bm{\mathbf{{U}}}}_{X}({\bm{\mathbf{{V}}}}_{X}{\bm{\mathbf{{v}}}})$ is cheaper to compute and approximates the product ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{v}}}}$ (for an arbitrary vector ${\bm{\mathbf{{v}}}}$). Computing matrix products with the data matrix ${\bm{\mathbf{{X}}}}$ is the main computational bottleneck for many NMF and SymNMF algorithms. This idea has been explored before in [48] where the authors used low-rank approximations such as the standard, truncated SVD and in [13] where the QB-decomposition was used.
@@ 3.1 SymNMF with Low-rank Approximate Input
We now present LAI-SymNMF. This algorithm is an instantiation of the low-rank approximate input method. Since the data matrix ${\bm{\mathbf{{X}}}}$ is symmetric we also require that our low-rank approximation be symmetric and not compress one side of the matrix ${\bm{\mathbf{{X}}}}$ (which would destroy symmetry) as in [13]. This is accomplished by using the approximate EVD of a symmetric matrix, using Algorithm Apx-EVD, which gives an approximate truncated EVD of ${\bm{\mathbf{{X}}}}$. The formulation for LAI-SymNMF is
$\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{\Lambda}}}}_{X}{\bm{\mathbf{{U}}}}_{X}^{\sf T}-{\bm{\mathbf{{H}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F},$ (21)
@@
2:${\bm{\mathbf{{H}}}}$ as the factors for an approximate rank-$k$ SymNMF of ${\bm{\mathbf{{X}}}}$.
@@
5:     [${\bm{\mathbf{{U}}}},{\bm{\mathbf{{\Lambda}}}}]:=$ Apx-EVD(${\bm{\mathbf{{X}}}},k,\rho,q$), $\triangleright$ Obtain approximate basis for range of ${\bm{\mathbf{{X}}}}$, ${\bm{\mathbf{{U}}}}\in\mathbb{R}^{m\times l}$ and ${\bm{\mathbf{{T}}}}\in\mathbb{R}^{l\times l}$
@@
9:         ${\bm{\mathbf{{X}}}}_{H}:=({\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{V}}}}){\bm{\mathbf{{U}}}}^{\sf T}+\alpha{\bm{\mathbf{{H}}}}^{\sf T}$triangleright This replaces ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{X}}}}$
@@
11:         ${\bm{\mathbf{{W}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{H}}}}},{\bm{\mathbf{{X}}}}_{H}\Big{)}$
12:         ${\bm{\mathbf{{X}}}}_{W}:=({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{U}}}}){\bm{\mathbf{{V}}}}^{\sf T}+\alpha{\bm{\mathbf{{W}}}}^{\sf T}$triangleright This replaces ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$
@@
14:         ${\bm{\mathbf{{H}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{W}}}}},{\bm{\mathbf{{X}}}}_{W}\Big{)}$
@@
LAI-SymNMF is flexible. Overall if the method for computing SymNMF requires computing the products ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$, and potentially ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ as in Equation 3, then LAI-SymNMF can be efficient. Not all methods for computing SymNMF rely on these products, such as Cyclic Coordinate Descent (CCD) [42]. As previously mentioned, in [25] it was shown that SymNMF via ANLS applied to Equation 3 is superior to PGD and a Newton-like method. Similarly, it was shown in [15] that SymNMF via BPP was competitive with a Projected Gauss-Newton based method for SymNMF and CCD. CCD is relatively inefficient for large problems as it iterates sequentially over every element of ${\bm{\mathbf{{H}}}}$. Therefore we use methods based on Equation 3 with BPP and HALS and the PGNCG’s method as our base line for comparison.
To simplify pseudocode and emphasize the flexibility of the method we introduce the $\text{Update}({\bm{\mathbf{{G}}}},{\bm{\mathbf{{Y}}}})$ function which takes in the Gram matrix ${\bm{\mathbf{{G}}}}\in\mathbb{R}^{k\times k}$ and the product between ${\bm{\mathbf{{X}}}}$ and either ${\bm{\mathbf{{W}}}}$ or ${\bm{\mathbf{{H}}}}$ denoted as ${\bm{\mathbf{{Y}}}}\in\mathbb{R}^{k\times m}$, and performs an update using methods such as HALS or BPP. For a more in-depth discussion of $\text{Update}()$ see Appendix G. We note that the Update() function abstraction is useful for Alternating Updating (AU) based methods. However, one of the advantages of our LAI method is that it is applicable to more algorithms. Existing randomized methods can be effectively used for the NMF update rules such as BPP, MU, or HALS but cannot or have not been used for all-at-once techniques such as PGNCG. Pseudocode showing how LAI can be used in conjunction with the PGNCG method is shown in Algorithm LAI-PGNCG-SymNMF in Section A.3.

### v2

@@ 2.1.3 Projected Gauss-Newton with Conjugate Gradients for SymNMF
The algorithm PGNCG-SymNMF was proposed for efficiently computing SymNMF in highly parallel computing environments [22]. It is an all-at-once method and uses the Projected Gauss-Newton method to directly optimize the SymNMF objective, Equation 2. The main computational load lies in solving a least squares problem of the form
@@
for a search direction ${\bm{\mathbf{{p}}}}$ at every iteration. The matrix ${\bm{\mathbf{{J}}}}$ and the vector ${\bm{\mathbf{{r}}}}$ are the Jacobian and residual of Equation 2 respectively. A solution to this LS problem is then approximated using the Conjugate Gradient (CG) method on the Normal Equations : ${\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{J}}}}{\bm{\mathbf{{p}}}}=-{\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{r}}}}$. A core computational kernel of the CG method is computing matrix vector products with the matrix ${\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{J}}}}$. Fortunately, the Jacobian ${\bm{\mathbf{{J}}}}$ has the Kronecker product form ${\bm{\mathbf{{J}}}}=-({\bm{\mathbf{{H}}}}\otimes{\bm{\mathbf{{I}}}}_{m})-{\bm{\mathbf{{P}}}}_{m,m}({\bm{\mathbf{{H}}}}\otimes{\bm{\mathbf{{I}}}}_{m})$, where ${\bm{\mathbf{{P}}}}_{m,m}$ is the perfect shuffle or “vec” permutation, which can be used to efficiently apply ${\bm{\mathbf{{J}}}}$ to a vector. Additionally the vector ${\bm{\mathbf{{g}}}}={\bm{\mathbf{{J}}}}^{\sf T}{\bm{\mathbf{{r}}}}$ has the form ${\bm{\mathbf{{g}}}}=-2\text{vec}({\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}-{\bm{\mathbf{{H}}}}({\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}))$, which is typically the main computational bottleneck and requires the matrix multiplications ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$. See [22] for details.
@@ 2.2.1 Randomized Range Finder
2:${\bm{\mathbf{{Q}}}}_{Y}\in\mathbb{R}^{m\times(r+\rho)}$, is an approximate orthonormal basis for the leading column span of ${\bm{\mathbf{{X}}}}$:
@@
An algorithm outline for the RRF is given in Algorithm RRF. Parameters of the RRF are the target rank $r$, a column oversampling parameter $\rho$, and $q$, the number of power iterations to perform. The computational complexity of the RRF is $O(qmnl+ml^{2})$ where $l=r+\rho$. The approximate output from the RRF is often used to compute a “QB-Decomposition”. If a matrix ${\bm{\mathbf{{X}}}}$ is input to the RRF and a matrix ${\bm{\mathbf{{Q}}}}_{X}$ is output, then ${\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{Q}}}}_{X}^{\sf T}{\bm{\mathbf{{X}}}}={\bm{\mathbf{{Q}}}}_{X}{\bm{\mathbf{{B}}}}_{X}$, where ${\bm{\mathbf{{B}}}}_{X}={\bm{\mathbf{{Q}}}}_{X}^{\sf T}{\bm{\mathbf{{X}}}}$, is called a QB-Decomposition of ${\bm{\mathbf{{X}}}}$.
@@
For a symmetric input, an approximate eigenvalue decomposition can also be obtained by using the approximate basis from the RRF. This procedure is shown in Algorithm Apx-EVD, which stands for approximate eigenvalue decomposition. More details and references can be found in [26].
@@ 2.2.2 Leverage Score Sampling for Ordinary Least Squares
Another approach that we will use from RandNLA is sketching for ordinary least squares (OLS) problems, specifically, leverage score sampling for OLS problems. A standard OLS or $l2-$regression problem is
(8) $\min_{{\bm{\mathbf{{x}}}}\in\mathbb{R}^{k}}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}$
@@
(9) $\hat{{\bm{\mathbf{{x}}}}}_{ols}=\operatorname*{arg\,min}_{{\bm{\mathbf{{x}}}}}\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{S}}}}{\bm{\mathbf{{b}}}}\|_{2},$
where ${\bm{\mathbf{{S}}}}\in\mathbb{R}^{s\times m}$, with $s\ll m$, is called a sketching matrix. Computational savings come from the fact that one can now solve the smaller problem in Equation 9 as opposed to the full sized problem in Equation 8.
@@
(10) $l_{i}({\bm{\mathbf{{A}}}})=\|{\bm{\mathbf{{Q}}}}_{A}[i,:]\|_{2}^{2}$
where the matrix ${\bm{\mathbf{{Q}}}}_{A}$ is any othonormal basis for the column space of ${\bm{\mathbf{{A}}}}$ and ${\bm{\mathbf{{Q}}}}_{A}[i,:]$ is the $i$th row of ${\bm{\mathbf{{Q}}}}_{A}$. For example the matrix ${\bm{\mathbf{{U}}}}_{A}$, where ${\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}^{\sf T}_{A}={\bm{\mathbf{{A}}}}$ is a thin SVD of ${\bm{\mathbf{{A}}}}$, can be used to calculate the leverage scores. These values are normalized into probabilities $p_{i}=\frac{l_{i}({\bm{\mathbf{{A}}}})}{\|{\bm{\mathbf{{Q}}}}_{A}\|_{F}^{2}}$. Using the $p_{i}$’s, $s$ samples are drawn with replacement and the matrix ${\bm{\mathbf{{S}}}}\in\mathbb{R}^{s\times m}$ is formed as
(11) ${\bm{\mathbf{{S}}}}_{ji}=\begin{cases}\frac{1}{\sqrt{sp_{i}}},\ \text{if row $i$ was drawn as the $j$th sample}\\ 0\end{cases}.$
@@
When leverage score sampling is used, with an appropriate number of samples $s$, the solution to the sampled problem in Eq. 9 satisfies the bound
(12) $\|\hat{{\bm{\mathbf{{x}}}}}_{ols}-{\bm{\mathbf{{x}}}}_{ols}\|_{2}\leq\sqrt{\epsilon_{r}}\frac{\|{\bm{\mathbf{{r}}}}_{ols}\|_{2}}{\sigma_{min}({\bm{\mathbf{{A}}}})}$
with probability $1-\delta$ where ${\bm{\mathbf{{r}}}}_{ols}={\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{ols}-{\bm{\mathbf{{b}}}}$ and ${\bm{\mathbf{{x}}}}_{ols}$ is the minimizer of Eq. 8. Additionally, $\delta$ and $\epsilon_{r}$ are values between 0 and 1 which control the failure probability and approximation error respectively. The number of samples $s$ is a function of these values and lower values of $\delta$ and $\epsilon_{r}$ incur a higher value of $s$. That is for lower failure probability and/or smaller error one must take more samples. Details can be found in [37].
2.2.3 Leverage Score Sampling for Nonnegative Least Squares
In order to apply leverage score sampling for Least Squares problems to NMF we extend results for Leverage Score Sampling for OLS problems to Nonnegative Least Squares Problems. The Nonnegative Least Squares problem has the general form
(13) $\min_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}.$
NLS problems are quite different from OLS problems as they do not yield a closed form solution. Despite this we are able to prove an analogous bound for sampling NLS problems with leverage scores. The statement of this result is given by the following theorem
Theorem 2.1.
Let ${\bm{\mathbf{{x}}}}_{nls}=\operatorname*{arg\,min}_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}$ be a NLS solution where ${\bm{\mathbf{{A}}}}\in\mathbb{R}^{m\times k}$, $m>k$, and $\text{rank}({\bm{\mathbf{{A}}}})=k$. Also let ${\bm{\mathbf{{S}}}}\in\mathbb{R}^{s\times m}$ be a leverage score sampling matrix for ${\bm{\mathbf{{A}}}}$ as in Eq. 11 with $s$ samples satisfying
for some $\epsilon_{r},\delta\in(0,1)$. Also let $\hat{{\bm{\mathbf{{x}}}}}_{nls}=\operatorname*{arg\,min}_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{S}}}}{\bm{\mathbf{{b}}}}\|_{2}$ be the sampled NLS solution. Then with probability $1-\delta$, the following holds :
where ${\bm{\mathbf{{r}}}}_{nls}={\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}}$ and $\sigma_{min}({\bm{\mathbf{{A}}}})$ is the minimum singular value of ${\bm{\mathbf{{A}}}}$ [^1] We note that Theorem 2.1 does not include a leverage score “missestimation factor” ($\beta$ in [56]) which is often included in works concerning leverage score sketching for OLS problems [37, 43, 56]. When inexact leverage scores are used for sampling, the missestimation factor gives a measure of how close the inexact leverage scores are to the true leverage scores. We do not thoroughly discuss missestimation factors because we do not use the concept in this work. However, Theorem 2.1 can be easily generalized to incorporate such a factor..
This result immediately begets the idea to use randomized NLS methods to rapidly solve the NLS subproblems in Alternating Nonnegative Least Squares methods for NMF and SymNMF previously described in Section 2.1.1. We described such a randomized algorithm in Section 4. A full proof of Theorem 2.1 is given in Section 4.3.1. We note that Theorem 2.1 holds for all convex Least Squares problems, this fact is made apparent in the proof given in Section 4.3.1.
2.3 Related Work
There has been much work on speeding up and scaling algorithms for low-rank methods mostly focusing on parallel algorithms and, more recently, randomization. We now give a brief review of related work with a focus on randomized methods.
2.3.1 Randomized Methods for other Low-rank Approximations
One of the first applications of leverage score sampling to matrix low-rank approximation was for computing a CUR decomposition [53]. The CUR decomposition selects actual rows and columns of the matrix to produce a low-rank approximation. Since the CUR decomposition is composed of sampled rows and columns it is said to be interpretable in the “space” of the original data.
There has been a lot of work related to randomized algorithms for computing the Canonical Polyadic (CP) decomposition for tensors. The CP Decomposition decomposes a $N$-way tensor into a sum of rank one tensors. Various constraints can be imposed similarly to low-rank matrix approximations, for example the Nonnegative CP decomposition. Two of the first methods for computing a randomized CP decomposition are by Battaglino et al. [1] and Zhou et al. [60]. Erichson et al. [18] proposed a randomized method for computing a CP decomposition based on using a tensor version of the RRF. The main idea is to first compress the tensor, via the RRF, compute a CP of the compressed tensor, and then lift the compressed CP back to the uncompressed space. Leverage scores have also been used in randomized CP algorithms [36, 6, 3]. There has also been work on scaling low-rank approximation methods to distributed computing environments [21, 29, 30] and combining such methods with randomized techniques [2]. These methods are applicable to sparse input tensors as they perform explicit sampling of tensor elements, thus preserving the sparsity pattern.
2.3.2 Randomized NMF Algorithms
Existing randomized algorithms for NMF focus on compressing the input matrix ${\bm{\mathbf{{X}}}}$. Speed up is obtained from the fact that iteratively updating the low-rank factors for a compressed version of ${\bm{\mathbf{{X}}}}$ is cheaper. The first proposed method we are aware of used a random Gaussian matrix to sketch, once from each side of ${\bm{\mathbf{{X}}}}$, resulting in two sketched matrices with smaller dimensions [54]. Tepper and Sapiro [51] proposed a similar method but used the RRF to compute an approximate basis for the row and column spans of ${\bm{\mathbf{{X}}}}$. Their results showed that using the RRF significantly increased the accuracy of the randomized NMF results in terms of final residual. Erichson et al. [19] proposed a randomized HALS algorithm also based on the RRF. This method computes a single RRF of the input matrix ${\bm{\mathbf{{X}}}}$ and fits NMF to the resulting QB-decomposition. This method has the advantage of calling the RRF once and avoids sketching the factor matrices of NMF at each iteration. However, due to the way the problem size is reduced, imposing nonnegativity to the factors of the original matrix becomes an issue.
@@ 3 NMF with Low-rank Approximate Input
Our first proposed algorithm is a method called Low-rank Approximate-Input NMF (LAI-SymNMF). LAI-NMF computes an NMF of a low-rank approximation of the initial data matrix ${\bm{\mathbf{{X}}}}$. The objective function for LAI-NMF is
(14) $\min_{\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}\geq 0}\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}-{\bm{\mathbf{{W}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F},$
where ${\bm{\mathbf{{U}}}}_{X}\in\mathbb{R}^{m\times l}$, ${\bm{\mathbf{{V}}}}_{X}\in\mathbb{R}^{l\times n}$ with $k\leq l\ll\min(m,n)$ and ${\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}\approx{\bm{\mathbf{{X}}}}$ is a low-rank approximation of ${\bm{\mathbf{{X}}}}$. The primary idea is that an approximate solution to Equation 14 can be quickly computed by exploiting the product form of ${\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}$ to compute matrix vector products. That is, ${\bm{\mathbf{{U}}}}_{X}({\bm{\mathbf{{V}}}}_{X}{\bm{\mathbf{{v}}}})$ is cheaper to compute and approximates the product ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{v}}}}$ (for an arbitrary vector ${\bm{\mathbf{{v}}}}$). Computing matrix products with the data matrix ${\bm{\mathbf{{X}}}}$ is the main computational bottleneck for many NMF and SymNMF algorithms. This idea has been explored before in [59] where the authors used low-rank approximations such as the truncated SVD, and in [19] where the QB-decomposition was used.
@@ 3.1 SymNMF with Low-rank Approximate Input
We now present LAI-SymNMF, an instantiation of the low-rank approximate input method. Since the data matrix ${\bm{\mathbf{{X}}}}$ is symmetric we also require that our low-rank approximation be symmetric and not compress only one side of the matrix ${\bm{\mathbf{{X}}}}$ (which would destroy symmetry) as in [19]. This is accomplished by using the approximate EVD of a symmetric matrix, using Algorithm Apx-EVD, which gives an approximate truncated EVD of ${\bm{\mathbf{{X}}}}$. The formulation for LAI-SymNMF is
(15) $\min_{{\bm{\mathbf{{H}}}}\geq 0}\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{\Lambda}}}}_{X}{\bm{\mathbf{{U}}}}_{X}^{\sf T}-{\bm{\mathbf{{H}}}}{\bm{\mathbf{{H}}}}^{\sf T}\|_{F},$
@@
2:${\bm{\mathbf{{H}}}}$, the factor for an approximate rank-$k$ SymNMF of ${\bm{\mathbf{{X}}}}$.
@@
5:     [${\bm{\mathbf{{U}}}},{\bm{\mathbf{{\Lambda}}}}]:=$ Apx-EVD(${\bm{\mathbf{{X}}}},k,\rho,q$), $\triangleright$ Obtain approximate, truncated EVD of ${\bm{\mathbf{{X}}}}$, ${\bm{\mathbf{{U}}}}\in\mathbb{R}^{m\times l}$ and ${\bm{\mathbf{{\Lambda}}}}\in\mathbb{R}^{l\times l}$
@@
9:         ${\bm{\mathbf{{Y}}}}_{H}:=({\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{V}}}}){\bm{\mathbf{{U}}}}^{\sf T}+\alpha{\bm{\mathbf{{H}}}}^{\sf T}$triangleright This replaces ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{X}}}}$
@@
11:         ${\bm{\mathbf{{W}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{H}}}}},{\bm{\mathbf{{Y}}}}_{H}\Big{)}$triangleright See Appendix E for description of Update()
12:         ${\bm{\mathbf{{Y}}}}_{W}:=({\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{U}}}}){\bm{\mathbf{{V}}}}^{\sf T}+\alpha{\bm{\mathbf{{W}}}}^{\sf T}$triangleright This replaces ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$
@@
14:         ${\bm{\mathbf{{H}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{W}}}}},{\bm{\mathbf{{Y}}}}_{W}\Big{)}$
@@
LAI-SymNMF is flexible. Overall if the method for computing SymNMF requires computing the products ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$ and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$, and potentially ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$ and ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$ as in Equation 3, then LAI-SymNMF can be efficient. Not all methods for computing SymNMF rely on these products, such as Cyclic Coordinate Descent (CCD) [52]. As previously mentioned, in [35] it was shown that SymNMF via ANLS applied to Equation 3 is superior to PGD and a Newton-like method. Similarly, it was shown in [22] that SymNMF via BPP was competitive with a Projected Gauss-Newton based method for SymNMF and CCD. CCD is relatively inefficient for large problems as it iterates sequentially over every element of ${\bm{\mathbf{{H}}}}$. Therefore we use methods based on Equation 3 with BPP and HALS, and the PGNCG’s method as our base line for comparison.
To simplify pseudocode and emphasize the flexibility of the method we introduce the $\text{Update}({\bm{\mathbf{{G}}}},{\bm{\mathbf{{Y}}}})$ function which takes in the Gram matrix ${\bm{\mathbf{{G}}}}\in\mathbb{R}^{k\times k}$ and the product between ${\bm{\mathbf{{X}}}}$ and either ${\bm{\mathbf{{W}}}}$ or ${\bm{\mathbf{{H}}}}$ denoted as ${\bm{\mathbf{{Y}}}}\in\mathbb{R}^{k\times m}$, and performs an update using methods such as HALS or BPP. For a more in-depth discussion of $\text{Update}()$ see Appendix E. We note that the Update() function abstraction is useful for Alternating Updating (AU) based methods. However, one of the advantages of our LAI method is that it is applicable to more algorithms. Existing randomized methods can be effectively used for the NMF update rules such as BPP, MU, or HALS but cannot or have not been used for all-at-once methods such as PGNCG. The algorithm outline showing how LAI can be used in conjunction with the PGNCG method is shown in Algorithm LAI-PGNCG-SymNMF in Section B.2.